跳到论文正文
W Papers独立论文阅读器 arXiv:1404.1161 论文正文 计算阅读时间 · 0%
引用这篇论文根据当前论文元数据生成,正式投稿前建议核对。

    
  
License: arXiv.org perpetual non-exclusive license
arXiv:1404.1161v1 [math.PR] 04 Apr 2014

The Probability Distribution for Draws Until First Success Without Replacement

John Ahlgren <<ahlgren@ee.cityu.edu.hk>>
Abstract

We consider the urn setting with two different objects, “good” and “bad”, and analyze the number of draws without replacement until a good object is picked. Although the expected number of draws for this setting is a standard textbook exercise, we compute the variance, and show that this distribution converges to the geometric distribution.

1 Introduction

Consider the setting where we draw from an urn containing two types of objects, “good” and “bad”. There are in total NN objects, of which KK are good.

Most instances of this setting are well known [1]: if we wish to obtain the number of good draws among nn many, use the binomial distribution for draws with replacement. Without replacement, we instead use the hypergeometric distribution. If, on the other hand, we wish to consider the number of draws until we draw a good object, we use the geometric distribution for draws with replacement. But what about draws without replacement? The situation is depicted in Table 1.

It is not too hard to show that the expected number of draws in this case is (N+1)/(K+1)(N+1)/(K+1), but what about the variance? Also, intuitively, such a distribution should be approximated by a geometric distribution if NN is much larger than KK, as the replacements will matter less.

Setting Counting Solutions Until First Solution
Replacements Binomial Geometric
No Replacements Hypergeometric ?
Table 1: Classification of Distributions.

In Section 2, we present the relevant probability mass function, cumulative function distribution, expected value, variance, and a convergence result. Longer proofs are given in the appendices. Section 3 concludes the paper.

2 The Probability Distribution

Consider an urn containing NN objects, of which KK are considered “good” and the remaining “bad”. The probability of picking a good object on the first trial n=1n=1 is clearly P⁡(1)=KNP(1)=\frac{K}{N}. For the second pick, we must have failed once and then succeeded: P⁡(2)=N−KN​KN−1P(2)=\frac{N-K}{N}\frac{K}{N-1}. To obtain the general form, we begin with an expression for the first nn failures.

Theorem 2.1.

The probability that our first nn picks are bad candidates (assuming independence) is

(N−K)​(N−K−1)​…​(N−K−n+1)N⁡(N−1)​…​(N−n+1)\frac{(N-K)(N-K-1)\ldots(N-K-n+1)}{N(N-1)\ldots(N-n+1)} (1)

which is well defined for n=1,2,…,(N−K)n=1,2,\ldots,(N-K).

Proof.

At step kk, k−1k-1 bad picks have been removed. Thus there are N−K−(k−1)N-K-(k-1) bad picks left and a total of N−(k−1)N-(k-1) picks left. The probability of failure is therefore (N−K−k+1)/(N−k+1)(N-K-k+1)/(N-k+1). There are nn failure steps, k=1,2,…,nk=1,2,\ldots,n, and since each step is independent, the result is obtained. Note that we can only fail at most N−KN-K times, as that is the total number of bad picks. ∎

By expressing Equation 1 as a ratio of binoms, we define the Fail-function.

Definition 2.2.

The Fail-function is defined as

F​a​i​l​(n)=(N−nK)(NK)Fail(n)=\frac{\binom{N-n}{K}}{\binom{N}{K}} (2)

for any integers KK, NN, and nn.

Due to Theorem 2.1, this corresponds to the probability of failure for n=1,2,…,(N−K)n=1,2,\ldots,(N-K). Thus the Fail-function extends expression 1 so that it is 00 when n>N−Kn>N-K and 11 for n=0n=0.

Since the probability of success at nn is simply n−1n-1 failures followed by a success, we can derive a closed form for the probability mass function.

Theorem 2.3 (Probability Mass Function).

The probability mass function (pmf) is:

P⁡(n)=F​a​i​l​(n−1)​KN−n+1=(N−n+1K)(NK)​KN−n+1.P(n)=Fail(n-1)\frac{K}{N-n+1}=\frac{\binom{N-n+1}{K}}{\binom{N}{K}}\frac{K}{N-n+1}. (3)
Proof.

We must fail n−1n-1 times and then succeed on our nnth time. Using Theorem 2.1 and noting that after having failed n−1n-1 times there are N−(n−1)=N−n+1N-(n-1)=N-n+1 candidates left, we obtain the expression for n≥2n\geq 2. For n=1n=1, Fail⁡(0)=1\mathrm{Fail}(0)=1 and we obtain the correct probability K/NK/N. ∎

To obtain a closed form for the cumulative distribution function, we take another approach.

Theorem 2.4 (Cumulative Distribution Function).

The cumulative distribution function (cdf) is:

F⁡(n)=∑k=1nP⁡(k)=1−F​a​i​l​(n)=1−(N−nK)(NK)F(n)=\sum_{k=1}^{n}P(k)=1-Fail(n)=1-\frac{\binom{N-n}{K}}{\binom{N}{K}} (4)
Proof.

The cmf is the probability of succeeding at any of the first nn picks, which is equivalent to not failing on any of the nn picks. ∎

Now that we have obtained the pmf and cdf, we are ready to present all results. For the remainder of this paper, let XX be a random variable with pmf as in Equation 3, that is, XX is the number of picks until we find a good enough solution.

Theorem 2.5 (Expected Value).
E⁡(X)=N+1K+1.\mathrm{E}(X)=\frac{N+1}{K+1}. (5)
Proof.

See Appendix A. ∎

The median does not have a nice closed form solution, but we can express it as a maximization problem using the cdf.

Theorem 2.6 (Median).

The median is the smallest integer mm such that

(N−mK)≤12​(NK).\binom{N-m}{K}\leq\frac{1}{2}\binom{N}{K}. (6)
Proof.

Immediate from Equation 4. ∎

The mode gives the most likely outcome.

Theorem 2.7 (Mode).

MODE⁡(X)=1\mathrm{MODE}(X)=1 if K>1K>1. If K=1K=1, P⁡(n)P(n) has fixed probability 1/N1/N for n=1,2,…,(N−K+1)n=1,2,\ldots,(N-K+1).

Proof.

See Appendix B. ∎

When K>1K>1, both picking with and without replacement are most likely to find a good solution in the first pick. Interestingly, for K=1K=1, we obtain a uniform distribution with P⁡(n)=1/NP(n)=1/N.

The variance gives the predictability of performance.

Theorem 2.8 (Variance).
VAR⁡(X)=K​(N−K)​(N+1)(K+2)​(K+1)2\mathrm{VAR}(X)=\frac{K(N-K)(N+1)}{(K+2)(K+1)^{2}} (7)
Proof.

See Appendix C. ∎

In the special case where K=1K=1, we have the variance for a uniform distribution, (N2−1)/12(N^{2}-1)/12.

Finally, we show that drawing without replacement can in some sense be approximated by its replacement counterpart as NN grows and the proportion of good solutions is held fixed.

Theorem 2.9.

Let K/NK/N be fixed. Let YY be a random variable modeling draws with replacement. X​⟶𝑑​YX\overset{d}{\longrightarrow}Y as N⟶∞N\longrightarrow\infty.

Proof.

See Appendix D. ∎

3 Conclusion

We have presented a probability distribution for number of draws until a first success is achieved among KK good and N−KN-K bad objects. Although the expected value of such a distribution is already known, we have also presented its variance and shown that the distribution converges towards a geometric distribution when NN grows and the proportion of good solutions K/NK/N is held fixed.

Appendix A Proof of Expected Value

First we start with a lemma.

Lemma A.1.
∑j=KN(jK)=(N+1K+1).\sum_{j=K}^{N}\binom{j}{K}=\binom{N+1}{K+1}.
Proof.

By induction over NN. For N=KN=K, the equality is trivial. Now assume it is valid for N=PN=P. Then, for N=P+1N=P+1, using Pascal’s triangle,

∑j=KP+1(jK)=(P+1K+1)+(P+1K)=(P+2K).\sum_{j=K}^{P+1}\binom{j}{K}=\binom{P+1}{K+1}+\binom{P+1}{K}=\binom{P+2}{K}. (8)

∎

Using Lemma A.1, we can obtain the expected value in closed form.

Proof of Expected Value.

First, reindex the sum and move out all factors independent of the summation index:

∑n=1N−K+1n​P​(n)=K(NK)​∑j=KNN+1−jj​(jK).\sum_{n=1}^{N-K+1}nP(n)=\frac{K}{\binom{N}{K}}\sum_{j=K}^{N}\frac{N+1-j}{j}\binom{j}{K}.

Expanding and using (ab)=ab​(a−1b−1)\binom{a}{b}=\frac{a}{b}\binom{a-1}{b-1}, we get two sums:

1(NK)​((N+1)​∑j=KN(j−1K−1)−K​∑j=KN(jK))\frac{1}{\binom{N}{K}}\left((N+1)\sum_{j=K}^{N}\binom{j-1}{K-1}-K\sum_{j=K}^{N}\binom{j}{K}\right)

which, by Lemma A.1, simplifies to

1(NK)​((N+1)​(NK)−K​(N+1K+1)).\frac{1}{\binom{N}{K}}\left((N+1)\binom{N}{K}-K\binom{N+1}{K+1}\right).

Simplifying, this yields the result. ∎

Appendix B Proof of Mode

To prove Theorem 2.7, we prove that the probability mass function PP is a decreasing function when K>1K>1 and constant when K=1K=1.

Theorem B.1 (Decreasing PMF).

Let n=1,2,…,(N−K+1)n=1,2,\ldots,(N-K+1). For K>1K>1, P⁡(n)>P⁡(n+1)P(n)>P(n+1). For K=1K=1, P⁡(n+1)=P⁡(n)=1/NP(n+1)=P(n)=1/N.

Proof.

Using Equation 3, for n=1,2,…,(N−K+1)n=1,2,\ldots,(N-K+1),

P⁡(n)P⁡(n+1)=N−nN−n+1−K\frac{P(n)}{P(n+1)}=\frac{N-n}{N-n+1-K} (9)

from which the result follows. ∎

Appendix C Proof of Variance

Before we prove the expression for variance, we will need two new lemmas. The first is a generalization of Lemma A.1.

Lemma C.1.

For x≥Kx\geq K,

∑j=xN(jK)=(N+1K+1)−(xK+1).\sum_{j=x}^{N}\binom{j}{K}=\binom{N+1}{K+1}-\binom{x}{K+1}.
Proof.

By writing

∑j=xN(jK)=∑j=KN(jK)−∑j=Kx−1(jK)\sum_{j=x}^{N}\binom{j}{K}=\sum_{j=K}^{N}\binom{j}{K}-\sum_{j=K}^{x-1}\binom{j}{K}

and then applying Lemma A.1 to both sums. ∎

Lemma C.2.
∑j=KNj​(jK)=N​(N+1K+1)−(N+1K+2).\sum_{j=K}^{N}j\binom{j}{K}=N\binom{N+1}{K+1}-\binom{N+1}{K+2}.
Proof.

The result is obtained by swapping rows and columns in the summation. First, consider the summation as a triangle:

∑j=KNj​(jK)=∑j=KN∑#​j(jK).\sum_{j=K}^{N}j\binom{j}{K}=\sum_{j=K}^{N}\sum_{\#j}\binom{j}{K}.

Transposing the sum, we get

K​∑j=KN(jK)+∑i=K+1N∑j=iN(iK).K\sum_{j=K}^{N}\binom{j}{K}+\sum_{i=K+1}^{N}\sum_{j=i}^{N}\binom{i}{K}.

The first sum can be solved using Lemma A.1, and the second using Lemma C.1, whereby we obtain

K​(N+1K+1)+∑i=K+1N((N+1K+1)−(iK+1)).K\binom{N+1}{K+1}+\sum_{i=K+1}^{N}\left(\binom{N+1}{K+1}-\binom{i}{K+1}\right).

Using Lemma A.1 once more, we obtain the result. ∎

Proof of Variance.

As with expected value, reindex the sum:

∑n=1N−K+1n2​P​(n)=1(NK)​∑j=KN(N+1−j)2​(j−1K−1)\sum_{n=1}^{N-K+1}n^{2}P(n)=\frac{1}{\binom{N}{K}}\sum_{j=K}^{N}(N+1-j)^{2}\binom{j-1}{K-1}

which after expanding the square and using (ab)=ab​(a−1b−1)\binom{a}{b}=\frac{a}{b}\binom{a-1}{b-1} gives:

1(NK)​((N+1)2​∑j=KN(j−1K−1)+K​∑j=KNj​(j−1K−1)CLOSE−2K(N+1)∑j=KNj(jK)).\frac{1}{\binom{N}{K}}\bigg((N+1)^{2}\sum_{j=K}^{N}\binom{j-1}{K-1}+K\sum_{j=K}^{N}j\binom{j-1}{K-1}\\ -2K(N+1)\sum_{j=K}^{N}j\binom{j}{K}\bigg). (10)

After using Lemma A.1 and C.2, and simplifying, we obtain the stated expression. ∎

Appendix D Proof of Convergence

Convergence of Distribution.

Set p=K/Np=K/N and q=(N−K)/Nq=(N-K)/N. Clearly K⟶∞K\longrightarrow\infty when N⟶∞N\longrightarrow\infty. Note that, for any nn, we have

N−K−n+1N−n+1=1−(1p+1−nK)−1⟶q​ as ​N⟶∞.\frac{N-K-n+1}{N-n+1}=1-\left(\frac{1}{p}+\frac{1-n}{K}\right)^{-1}\longrightarrow q\text{ as }N\longrightarrow\infty.

This implies that

F​a​i​l​(n)=∏k=1nN−K−k+1N−k+1⟶qnFail(n)=\prod_{k=1}^{n}\frac{N-K-k+1}{N-k+1}\longrightarrow q^{n}

so that

P⁡(n)=F​a​i​l​(n−1)​(1−(N−K−n+1N−n+1))⟶qn−1​(1−q)​ as ​N⟶∞.P(n)=Fail(n-1)\left(1-\left(\frac{N-K-n+1}{N-n+1}\right)\right)\\ \longrightarrow q^{n-1}(1-q)\text{ as }N\longrightarrow\infty. (11)

∎

Acknowledgements

The work described in this paper was supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China [Project No. CityU 125313].

References

  • [1] E.T. Jaynes and G.L. Bretthorst. Probability Theory: The Logic of Science. Cambridge University Press, 2003.